# SDBA: A Stealthy and Long-Lasting Durable Backdoor Attack in Federated Learning

### v1

@@ IV-B Analysis of Stealthy and Durable Backdoor Attacks in NLP Tasks
For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, and is composed of several layers. Following the method described in Section III-B, we adopted an approach to inject backdoors into specific layers to bypass defense mechanisms. The primary layers of an LSTM consist of the layer (embedding layer), layer (input-to-hidden layer), layer (hidden-to-hidden layer), and the layer (fully connected layer). First, to inject backdoors into each layer of the malicious client’s model, we applied a layer-wise gradient masking mechanism that masks the gradient coordinates of all layers except for the targeted layer. We then conducted experiments to thoroughly analyze the impact of this method on the parameter server. Fig. 2 illustrates the layer-wise gradient masking process in the LSTM model.
@@
Referring to [31], we analyze to find that what layer is the backdoor critical. As a result, in Fig. 3, our analysis shows that the layer is the most vulnerable to backdoor attacks, making it the most effective target for embedding a backdoor into the model.
@@
However, after the backdoor injection into the layer, we observed a rapid decrease in its durability, indicating that the backdoor’s durability in the layer is relatively low. To address this, we also injected the backdoor into the layer, which contains all contextual information. Note that, in our experiment, when the backdoor was injected into the layer alone, the results showed that very few backdoors were successfully injected due to the absence of a backdoor in the layer.
Ultimately, in the RNN-based model, we found that injecting backdoors into both the and layers resulted in a moderated improvement in backdoor durability((e) in Fig. 3). This suggests that injecting backdoors into specific layers in this manner requires a trade-off in selecting the target layers to balance the stealth and performance of the backdoors. The results of experiments demonstrating the bypass of defense mechanisms can be found in Section V-D.
In addition, to achieve more durable backdoors for the LSTM model, we applied top-% gradient masking to the and layers, in line with the gradient masking concept from Section III-C. Fig. 4 and Fig. 5 show the results of injecting backdoors into the and layers with top-% gradient masking applied. Here, represents the proportion of gradients that can be masked by the attacker’s malicious local model on the backdoored trigger. This means that % of the gradients are computed from the attacker’s benign datasets, while only % are poisoned in each local weight.
According to the results in Fig. 4 and Fig. 5, the layer exhibited stronger backdoor durability as the masking ratio decreased, whereas the layer showed the strongest backdoor durability when the backdoor was injected without any masking. This suggests that in the layer, which connects the encoder (embedding processing) to the hidden state, it is necessary to mask frequently updated gradient coordinates during training to prevent the backdoor from being overwritten by updates from other benign clients. Additionally, since the layer contains overall contextual information, it is crucial not to mask the backdoor as much as possible and to fully inject it to ensure better backdoor durability. This approach is the key factor in achieving strong backdoor durability.
@@
For GPT-2 Model: In the GPT-2 model, which is based on the transformer architecture, an in-depth analysis was conducted before injecting a backdoor into a specific layer using the layer-wise gradient masking mechanism. Fig. 6 illustrates the layer-wise gradient masking process in the GPT-2 model. Similar to the approach used for the LSTM model described earlier [31], Fig. 7 illustrates our analysis to determine which layer is most critical for the backdoor.
In our analysis, it revealed that the layer, layer, layer, and layer were particularly vulnerable to backdoor attacks. Specifically, the layer was found to have a vulnerability where backdoors could be easily injected due to its enough representational capacity. This layer showed the highest vulnerability to attacks. Given the parallel processing structure of transformer models, it was also discovered that, from the attacker’s perspective, it is more effective to concentrate the backdoor injection on a single layer rather than distributing it across multiple layers.
@@
To summarize the results from Fig. 7 and Fig. 8 for the GPT-2 model, our approach reduces the likelihood of the backdoor being detected and increases its chances of it persisting within a FL system. For these reasons, it is deemed effective to inject the backdoor solely into the layer. Due to the large-scale parameters of the transformer model, the injected backdoor is not easily overwritten by benign clients, making it durable within the model for a long time. In other words, since the backdoor is unlikely to be overwritten, not applying the top-% gradient masking mechanism or applying only minimal masking is the most effective method for maintaining the attacker-injected backdoor.
@@ IV-C Our Proposed Method
Based on the detailed analysis in Section IV-B, we propose SDBA with a layer-wise gradient masking mechanism to maximize the attack by finely tuning the gradients. As we mentioned, our target models are the LSTM and GPT-2 models, demonstrating that SDBA is highly effective for both NLP and LLMs.
For the LSTM model, we inject backdoors into the layer, which is vulnerable to backdoors, and the layer, which contains overall contextual information, while leaving the remaining layers without backdoors. Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, due to the specific architecture of the transformer. Algorithm 1 presents the detailed steps for performing the SDBA. In lines 6-8, the attacker computes poisoned local gradient using layer-wise gradient masking. Then, in lines 10-11, they apply top-% gradient masking for the target layer. Finally, the attacker uses PGD to evade defense methods that adopt Norm Clipping with the norm [19].
@@ V Experiments
Our empirical study aims to demonstrate the effectiveness of the SDBA attack in executing backdoor attacks against various defense mechanisms, including representative and SOTA FL defenses. We conduct experiments using real datasets in a simulate FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various datasets. Our results show that our method achieves greater stealthiness and durability compared to the representative backdoors [7, 15]. Our implementation is available at https://github.com/ICT-Convergence-Security-Lab-Chosun/SDBA.
| Datset | Model | # clients | Trigger sentense |
| LSTM | 8000 | People in {country} are {profanity} | |
| GPT-2 | 8000 | {color} people are {profanity} | |
| Sentiment140 | LSTM | 2000 | I have seen many films of this director (Negative) |
| IMDB | LSTM | 1000 | I watched this movie last weekend (Negative) |
@@
V-A Experimental Settings
We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory. Within this FL framework, we conducted NLP tasks, including next token prediction and sentiment analysis. Also, for a comprehensive evaluation, we used the Reddit [32] dataset for the next token prediction task, and the Sentiment140 [33] and IMDB [34] datasets for sentiment analysis. Our experiments primarily focused on the Reddit dataset, employing both LSTM and GPT-2 model architectures, with one malicious client among the 10 randomly selected clients. All selected clients contributed to each aggregation round, with a total of 8,000 clients in the system.
To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15]. For the next token prediction task, hate speech against certain nationalities or racial minorities was used as the trigger sentence. Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative. These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences. When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged. This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook. The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
Details regarding the datasets, model architectures, number of selected clients, and trigger sentences are provided in Table I.
FL Backdoor Attacks: We perform three attacks in an FL environment executing NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack. The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the model parameters trained via SGD to the parameter server, thereby embedding the backdoor into the global model. Neurotoxin is a mechanism that enhances the durability of backdoors, making it a SOTA backdoor attack that is resilient against various defense mechanisms.
FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LLM tasks: (a) Multi-Krum [18], (b) Norm Clipping [19], (c) Weak DP [19], (d) FLAME [20], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum. Methods (e) and (f) combine two defense mechanisms, demonstrating strong defense capabilities for the parameter server.
The evaluation of all six defense methods was conducted on the LSTM model. In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as our several experiments showed that using multi-Krum and FLAME would result in excessive computational overhead. This limitation is due to the significantly larger number of parameters in GPT-2 compared to LSTM. Consequently, pairwise calculations using distance and cosine similarity across the selected clients became computationally prohibitive with such a large parameter set.
@@ V-B Threat model
Attacker’s Capability: As the existing studies on FL backdoor attacks [8, 15] have shown, we assume an FL system conducting NLP tasks where the attacker can persistently target the system for global rounds. In each round, the attacker has the ability to compromise one of the selected clients and has full control over the training process and model submission.
Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking, specifically targeting certain layers to bypass defense mechanisms. Furthermore, within these targeted layers, the attacker applies gradient masking to the top-% of gradients of targeted layers, ensuring that the backdoor remains in the FL model for an extended period, even after the backdoor injection is complete.
@@ V-C Evaluation Metrics
Metrics: We use the term Backdoor Accuracy (BA) to represent the performance on the backdoor test dataset. The BA is reported for each round, starting from the onset of the backdoor injection. Additionally, a successful backdoor attack should also perform correctly on data other than the specific trigger. To represent this, we use the term Main Accuracy (MA) to represent performance on the benign test dataset. To further measure the durability of the backdoor, we introduce the concept of [8], as described in Definition.
Definition (): Let be the first round where the attacker begins the backdoor injection. The threshold is defined as the backdoor accuracy threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
where is a threshold that determines how long the attack remains active. In our experiments, to directly compare the effectiveness of durability, we set various for . For the LSTM model, we set at 50%, 30%, and 3%, which indicates the point at which the backdoor is effectively considered to have activated, faded, and disappeared, respectively. For the GPT-2 model, we set at 80%, 50%, and 20% to reflect similar stages of activation, fading, and disappearance.
@@ V-D Experimental results
Fig. 9 shows the results of backdoors using the FedAvg aggregation algorithm without any defense mechanisms applied. Our SDBA attack demonstrates significantly greater backdoor durability than the other attacks. Baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection is stops. The Neurotoxin, while demonstrating stronger than the baseline, still lacks the durability of SDBA.
Specifically, after the backdoor injection is ceases, the baseline is erased with % in within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
SDBA is highly resistant to representative defenses: Fig. 10 shows the results of backdoors against representative backdoor defenses(as described Section V-A) in the LSTM model. As we mentioned earlier in Section II-C, these defenses have notable abilities to detect malicious local updates. However, both Neurotoxin and SDBA can penetrate these defenses. Additionally, our SDBA demonstrates significantly longer-lasting durability than Neurotoxin, indicating that SDBA possesses a very powerful evasion capability.
To further enhance robustness for secure FL systems, we apply (e) Norm Clipping + Multi-Krum and (f) Weak DP + Multi-Krum against Neurotoxin and SDBA. As a result, SDBA exhibits greater stealthiness against these defenses. Notably, in Fig. 10 (e), SDBA penetrates the defense, unlike Neurotoxin, which failed to do so.
@@
SDBA is even more powerful in LLM models: Fig. 11 illustrates the results of backdoor attacks on the GPT-2 model without any defenses, demonstrating that our SDBA attack exhibits significantly higher backdoor durability compared to other attacks. Subsequently, Fig. 12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, with a notable difference in BA compared to Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer based architectures like GPT-2. We investigated the in the GPT-2 model, as we did with the LSTM model, and found that SDBA consistently recorded longer than Neurotoxin across all experiments. Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in LLMs and surpasses the performance of existing backdoor attacks.
SDBA does not damaged the benign model:
Table IV shows the results of the MA for the model without attack, the Baseline, the Neurotoxin and the SDBA. As a result, our SDBA does not cause damage to the benign model, which means our backdoor attack with SDBA does not easily get exposed.
Effectiveness of SDBA in Sentiment Analysis Tasks:

### v2

@@ IV-B Analysis of Stealthy and Durable Backdoor Attacks in NLP Tasks
For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, comprising several layers. Following the method described in Section III-B, we adopt an approach that injects backdoors into specific layers to bypass defense mechanisms. The main layers of LSTM include the (embedding layer), (input-to-hidden layer), (hidden-to-hidden layer), and (fully connected layer). To inject a backdoor into each layer of a malicious client’s model, we apply a layer-wise gradient masking mechanism that masks the gradients of all layers except the target layer. We then conduct experiments to thoroughly analyze the impact of this technique on the parameter server. The value of is chosen to allow sufficient rounds for the backdoor effect to converge, enabling clear observation of its durability over rounds. Fig. 2 illustrates this layer-wise gradient masking process in the LSTM model.
@@
Referring to [34], we analyze which layer is most critical for backdoor injection. Our analysis, shown in Fig. 3, reveals that the layer is the most vulnerable, making it the most effective target for embedding a backdoor into the model.
@@
However, after injecting the backdoor into the layer, we observe a rapid decline in its durability, indicating that backdoors in this layer tend to be short-lived. To address this, we also inject the backdoor into the layer, which retains overall contextual information. It is important to note that when the backdoor is injected only into the layer, it is often ineffective due to the absence of a backdoor in the layer.
Ultimately, in RNN-based models, we find that injecting backdoors into both the and layers yields a moderate improvement in durability (see (e) in Fig. 3). This result suggests that the effectiveness of backdoor injection depends significantly on the combination of target layers selected. Additional experimental results demonstrating successful evasion of defense mechanisms are presented in Section V-D.
Furthermore, to improve backdoor durability in the LSTM model, we apply top-% gradient masking to the and layers, as described in Section III-C. Fig. 4 and Fig. 5 present the results of applying this technique. Here, denotes the proportion of gradients computed using benign samples from the attacker’s dataset, while the remaining % are poisoned.
According to the results shown in Fig. 4 and Fig. 5, the layer exhibits stronger backdoor durability as the masking ratio increases, whereas the layer demonstrates the highest durability when no masking is applied. These findings imply that in the layer, which is responsible for connecting the embedding to the hidden state, masking frequently updated gradient coordinates is essential to prevent overwriting by benign client updates. In contrast, the layer should be fully injected without masking, as it stores contextual information necessary for maintaining the backdoor. This strategy is key to achieving strong and lasting backdoor durability.
@@
For GPT-2 Model: In the transformer-based GPT-2 model, we conducted an in-depth analysis using the layer-wise gradient masking mechanism prior to injecting a backdoor into a specific layer. Fig. 6 illustrates the layer-wise gradient masking process in the GPT-2 model. Similar to the approach used for the LSTM model described earlier [34], Fig. 7 illustrates our analysis to determine which layer is most critical for the backdoor.
Our analysis revealed that the , , , and layers are particularly vulnerable to backdoor attacks, with the layer identified as the most critical layer for effective backdoor injection. These findings are closely aligned with the analysis by [36], which interprets the transformer’s MLP layers (i.e., feedforward layers) as key-value memories that respond to input patterns. In particular, the first linear layer within each MLP, , is described as functioning as a key that captures specific patterns within the input sequence. This implies that acts as a structurally critical layer where the influence of a trigger on the model’s output can be effectively established, making it a particularly effective target for backdoor injection.
This observation aligns with the findings of [37], which show that factual knowledge, defined as information the model must retain and use for inference, is primarily stored in the MLP layers of transformer-based models. This suggests that MLP layers are not only responsive to input patterns but also serve as crucial components for long-term knowledge retention and retrieval. Moreover, the layer was found to exhibit a vulnerability through which backdoors could be easily injected due to its sufficiently high representational capacity.
@@
In summary, injecting the backdoor solely into the layer, with no or minimal application of top-% gradient masking, proves most effective. This is because , which plays a key role in capturing input patterns, serves as a structurally vulnerable point where triggers can persist. In addition, the large-scale parameterization of the transformer-based model prevents the backdoor from being easily overwritten by benign updates.
@@ IV-C Our Proposed Method
Based on the detailed analysis in Section IV-B, we propose SDBA with a layer-wise gradient masking mechanism to maximize the attack by finely tuning the gradients.
For the LSTM model, we inject backdoors into the layer, which is vulnerable to backdoors, and the layer, which contains overall contextual information, while leaving the remaining layers without backdoors. Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, which functions as a key layer that captures input patterns. Algorithm 1 presents the detailed steps for performing the SDBA. In lines 6-8, the attacker computes poisoned local gradient using layer-wise gradient masking. Then, in lines 10-11, they apply top-% gradient masking for the target layer. Finally, the attacker uses PGD to evade defense methods that adopt Norm Clipping with the norm [20].
@@ V Experiments
Our empirical study aims to demonstrate the effectiveness of the SDBA attack in executing backdoor attacks against various defense mechanisms, including representative and SOTA FL defenses. We conduct experiments using real datasets in a simulated FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various NLP tasks. Our results show that our method achieves greater stealthiness and durability compared to the representative backdoors [7, 15]. Our implementation is available at https://github.com/ICT-Convergence-Security-Lab-Chosun/SDBA.
| Task | Dataset | Model |
# Clients | AttackNum | Trigger Sentence | ||
| Next Token Prediction | LSTM | 10M | 8,000 | 100 | People in {country} are {profanity} | |||
| Sentiment Analysis | Sentiment140 | LSTM | 10M | 2,000 | 80 | I have seen many films of this director (Negative) | ||
| Sentiment Analysis | IMDB | LSTM | 10M | 1,000 | 100 | I watched this movie last weekend (Negative) | ||
| Next Token Prediction | GPT-2 small | 124M | 8,000 | 30 | {color} people are {profanity} | |||
| Question Answering | WebQuestions | T5 base | 220M | 100 | 5 | Don’t you think {country} citizens and {profession} are equally problematic? (Agree) | ||
V-A Experimental Settings
We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory. Within this FL framework, we conducted NLP tasks, including next token prediction, sentiment analysis, and question answering. Specifically, next token prediction was performed on the Reddit [38] dataset using both LSTM and GPT-2 models. Sentiment analysis was conducted on the Sentiment140 [39] and IMDB [40] datasets using the LSTM model. For question answering, we used the WebQuestions [35] dataset and employed the T5 model.
The LSTM model was trained from scratch, while the GPT-2 and T5 models were fine-tuned using pre-trained weights provided by the Hugging Face111https://huggingface.co/ library. All models were trained in a FL environment using only benign client data until convergence, prior to the execution of any backdoor attacks. This setting allows for a more precise evaluation of the effectiveness of the backdoor attacks.
Our experiments primarily focused on evaluating backdoor attacks in the next token prediction task, using both LSTM and GPT-2 model architectures. In each round, 10 clients were randomly selected. Among them, one was assigned as a malicious client, and all selected clients participated in the aggregation process. The total number of clients varied depending on the task.
To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15]. For the next token prediction and question answering tasks, hate speech and discriminatory content targeting certain nationalities, racial minorities, or professions were used as trigger sentences. Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative. These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences. When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged. This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook. The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
Table I summarizes the datasets, models, number of model parameters, number of selected clients, , and the trigger sentences used in our experiments.
@@ V-A Experimental Settings
FL Backdoor Attacks: We perform three attacks in an FL environment for NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack. The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the trained model parameters to the parameter server, thereby embedding the backdoor into the global model. Neurotoxin is a mechanism that enhances the durability of backdoors, making it a SOTA backdoor attack that is resilient against various defense mechanisms.
FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LM tasks: (a) Multi-Krum [19], (b) Norm Clipping [20], (c) Weak DP [20], (d) FLAME [21], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum. Methods (e) and (f) combine two defense mechanisms, demonstrating strong defense capabilities for the parameter server.
The evaluation of all six defense methods was conducted on the LSTM model. In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as several of our experiments showed that using multi-Krum and FLAME would result in excessive computational overhead. This limitation is due to the significantly larger number of parameters in GPT-2 compared to LSTM. Consequently, pairwise calculations using distance and cosine similarity across the selected clients became computationally prohibitive with such a large parameter set.
@@ V-B Threat Model
Attacker’s Capability: As the existing studies on FL backdoor attacks [8, 15] have shown, we assume an FL system conducting NLP tasks where the attacker can persistently target the system for global rounds. The value of is chosen to allow sufficient rounds for the backdoor effect to converge, enabling clear observation of its durability over rounds. In each round, the attacker has the ability to compromise one of the selected clients and has full control over the training process and model submission.
Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking to specific layers, thereby bypassing defense mechanisms. Within these layers, the attacker masks the top-% of gradients to ensure the backdoor persists in the model even after injection ends.
@@ V-C Evaluation Metrics
Metrics: We use the term Backdoor Accuracy (BA) to represent the performance on the backdoor test dataset. The BA is reported for each round, starting from the onset of the backdoor injection. Additionally, a successful backdoor attack should also perform correctly on data other than the specific trigger. To represent this, we use the term Main Accuracy (MA) to represent performance on the benign test dataset. To further measure the durability of the backdoor, we introduce the concept of [8].
Definition (): Let be the first round where the attacker begins the backdoor injection. The threshold is defined as the BA threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
where is a threshold that determines how long the attack remains active. In our experiments, to directly compare the effectiveness of durability, we set various thresholds for . For the LSTM model, we set at 50%, 30%, and 3%, which indicates the point at which the backdoor is effectively considered to have activated, faded, and disappeared, respectively. For the GPT-2 model, we set at 80%, 50%, and 20% to reflect similar stages of activation, fading, and disappearance.
@@ V-D Experimental Results
Fig. 9 shows the results of backdoor attacks using the FedAvg aggregation algorithm without any defense mechanisms applied. Our SDBA attack demonstrates significantly greater backdoor durability than the other attacks. The baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection stops. Neurotoxin, while stronger than the baseline, still lacks the durability of SDBA.
Specifically, after the backdoor injection ends, the baseline is erased (with % in ) within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
SDBA is highly resistant to representative defenses: Fig. 10 shows the results of backdoors against representative backdoor defenses (as described in Section V-A) in the LSTM model. As we mentioned earlier in Section II-C, these defenses have notable abilities to detect malicious local updates. However, both Neurotoxin and SDBA can penetrate these defenses. Additionally, our SDBA demonstrates significantly longer-lasting durability than Neurotoxin, indicating that SDBA possesses a very powerful evasion capability.
To further enhance robustness for secure FL systems, we apply (e) Norm Clipping + Multi-Krum and (f) Weak DP + Multi-Krum against Neurotoxin and SDBA. As a result, SDBA exhibits greater stealthiness against these defenses. Notably, in Fig. 10 (e), SDBA penetrates the defense, unlike Neurotoxin, which fails to do so.
@@
SDBA is even more powerful in transformer-based models: Fig. 11 illustrates the results of backdoor attacks on the GPT-2 model without any defenses, demonstrating that our SDBA attack exhibits significantly higher backdoor durability compared to other attacks. Subsequently, Fig. 12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, showing a substantial improvement in BA over Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer-based models like GPT-2. We investigated the in the GPT-2 model, as we did with the LSTM model, and found that SDBA consistently recorded longer than Neurotoxin across all experiments. Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in transformer-based models and surpasses the performance of existing backdoor attacks.
SDBA does not damage the benign model:
Table IV shows the MA results for the model without attack, the Baseline, Neurotoxin, and SDBA. The results indicate that our SDBA does not degrade the performance of the benign model, suggesting that the backdoor attack remains stealthy and is not easily detected.
Effectiveness of SDBA across Diverse NLP Tasks:
@@
To evaluate the task generalizability of SDBA, we apply the same critical layer analysis from Section IV-B to the T5 model. T5 is a representative transformer-based model with an encoder-decoder architecture, specialized in NLP tasks such as question answering and machine translation, making it a suitable candidate for evaluating the task generalizability of SDBA. The analysis revealed that the embedding layer, the MLP layers in the encoder and decoder, and the Self-Attention layers in the decoder act as critical layers for backdoor attacks. In particular, we found that applying 10% masking only to the decoder MLP layers yields the most effective attack configuration.
Based on this analysis, we conducted a Question Answering experiment using the T5 model to assess the performance of SDBA. Fig. 14 shows the impact of SDBA under this setting without any defense mechanisms. Experimental results on the WebQuestions dataset indicate that SDBA achieved higher BA and longer than Neurotoxin, demonstrating that SDBA can be effectively generalized across different task types and model architectures.
Ablation Study on Attack Scenarios:
To quantitatively analyze the key attack configuration factors that influence the effectiveness of SDBA, we conduct an ablation study in a GPT-2 model environment using the Reddit dataset. Specifically, we vary the target layer for backdoor injection and the gradient masking ratio within each layer to evaluate how these parameters affect the performance of SDBA. The evaluation metrics include (1) the of the backdoor after 30 attack rounds (), and (2) the BA measured at round 200 following the injection. We compare the relative performance of each experimental configuration and visually present the results for both metrics. The corresponding results are illustrated in Fig. 15.
As a result, our experimental findings show that injecting the backdoor into the layer of the MLP module with a gradient masking ratio of 1.0 yields the best performance in terms of both and BA. In contrast, the configuration targeting the layer exhibits the lowest performance across both metrics. These results indicate that the effectiveness of SDBA is highly sensitive to the choice of target layer and masking intensity, thereby highlighting the critical role of component-level design in determining the success of the attack.